OpenStudy (anonymous):

x-2y=6 (use a table of values to graph the eqation) please help me with this graphing problem

OpenStudy (anonymous):

x - 2y = 6 -2y = -x + 6 y = -x/-2 + 6/-2 y = x/2 - 3 x = 0 , y = -3 x = 2, y = -2 x = 4, y = -1 Plot these three ordered pairs and draw a line through these points.